OpenStudy (anonymous):

Simplify (1 − cos x)(1 + cos x). 1 cos2 x sin2 x tan2 x @paki @ilikemath50 @FriedRice

OpenStudy (abmon98):

1+cosx-cosx-cos^2x

OpenStudy (abmon98):

1-cos^2x 1-(1-sin^2x)=sin^2x

OpenStudy (anonymous):

@Abmon98 what are you getting at

OpenStudy (anonymous):

@marissalovescats can you help

OpenStudy (marissalovescats):

Well when you foil (1 − cos x)(1 + cos x) you get 1-sin^2x And if you look on a trig identities sheet you'll see that sin^2x+cos^2x=1 and you subtract cos^2x from each side which tells you that sin^2x=1-cos^2x. So we know the answer to this is sin^2x because when you foil (1 − cos x)(1 + cos x) you get 1-cos^2x and 1-cos^2x=sin^2x.

OpenStudy (marissalovescats):

I meant you get 1-cos^2x in the first line sorry